OpenStudy (anonymous):

if z^3xy+4xz-sinxy=2xy+4yz^2 implicitly defines y as a function of x and z, find Fy/Fz

OpenStudy (anonymous):

\[z^3xy+4xz-(\sin x)y=2xy+4yz^2\] or\[z^3xy+4xz-\sin (xy)=2xy+4yz^2~~?\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

One more question, do you mean \(\dfrac{F_y}{F_z}\) to mean the ratio of the function's partial derivatives? or the derivative of y with respect to z?

OpenStudy (anonymous):



OpenStudy (anonymous):

derivative of y with respect to z

OpenStudy (anonymous):

Do you still need help with this?

OpenStudy (anonymous):

no. I figured it out! Thanks though!